OpenStudy (anonymous):

Which of the following equations is equivalent to -2(-3 + 2x) = 5? A. 5 = 5 B. 6 − 4x = 5 C. 6 − 4x = -10 D. 6 − 4x = 10

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Umm, use distributive property. \[ -2(-3 + 2x) = -2\cdot-3 + -2\cdot2x \]

OpenStudy (anonymous):

distribute the -2 and then combine like terms.

OpenStudy (anonymous):

-2(-3+2x)=5 (-2)(-3)+(-2)(2x)= 5 6-4x=5 :)

OpenStudy (anonymous):

That would be letter B ^.^ hope it helped..